A ball falls from rest under gravity on the moon such that 1st 1/2 of distance is vovered in 10seconds toatl time of the fallof ball is
 

Grade:11

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

S= 0.5 (a t*t)

H/2 = 0.5 ( a*100)

Using H = 100a

Total time for 100a from start
extra time required = 10((sqrt2)-1)

Regards

100a = 0.5a T *T
T = 10(sqrt(2))